# Calibrating Bayesian Tension Statistics using Neural Ratio Estimators

### v1

@@ VII.1 Toy 21-cm Cosmology
Observers in the field of 21-cm Cosmology are aiming to detect an information rich redshifted signal from neutral hydrogen from the Cosmic Dawn and Epoch of Reionization (see Furlanetto et al., 2006; Barkana, 2016; Mesinger, 2019; Liu and Shaw, 2020, for reviews of the field). The signal is observed in the radio band, and can in theory be detected as a sky-averaged 21-cm signal (e.g. Bowman et al., 2018; Singh et al., 2022b; de Lera Acedo et al., 2022). It has a complex dependence on the astrophysics of the early Universe (e.g. Mirocha, 2014; Mesinger et al., 2011; Reis et al., 2020, 2021; Gessey-Jones et al., 2022; Sikder et al., 2024; Pochinda et al., 2024; Gessey-Jones et al., 2024; Muñoz, 2023), but it can be approximated by a Gaussian absorption feature (e.g. Anstey et al., 2021) in the CMB spectrum akin to a spectra distortion. The key challenge in 21-cm cosmology is the separation of this signal from the dominant Galactic and extragalactic foregrounds, that the instruments also observe, whilst accounting for the non-uniform response of the instruments to the sky (Anstey et al., 2021).
We ignore the effects of foregrounds and the instrument in our example, since we are focused on illustrating the performance of the tensionnet. We include Gaussian distributed noise in our simulated data (inspired by current observations Bowman et al. (2018); Singh et al. (2022b)) and a Gaussian absorption feature
@@
where $A$ corresponds to the amplitude of the signal, $\nu_{0}$ to the central frequency and $w$ to the width.
@@
Current observations of the sky-averaged 21-cm signal include a tentative detection by the EDGES collaboration Bowman et al. (2018) and an upper limit on the magnitude of the signal from SARAS3 Singh et al. (2022b). Analysis by the SARAS3 team suggested that these measurements are in tension with each other, and a number of works have discussed the possible presence of systematics in the EDGES data Hills et al. (2018); Singh and Subrahmanyan (2019); Sims and Pober (2020); Bevins et al. (2021). As more experiments come online in the coming years (e.g. de Lera Acedo et al., 2022) the assessment of tension and concordance between different observations is going to become crucial for the field.
@@
Figure 5: To further illustrate the application of NREs to the calibration of $R$ we use a toy example inspired by 21-cm cosmology. Left Panel: We simulate an experiment observing a Gaussian absorption trough as a function of frequency (black line) and three different scenarios in which another experiment measures a 21-cm signal with either the same or different amplitudes in a different band (red lines). To each observation, we add Gaussian random noise with a standard deviation of 25 mK (shown in grey and motivated by current observations (e.g. Bowman et al., 2018)). Middle Panel: We train the NRE on simulated observations of the signal by both experiments, covering a wide prior range of signal parameters. We use the NRE to evaluate the possible distribution of in concordance $\log R$ values, which is shown in the middle panel. We plot the observed $\log R$ for each pair of observations from experiment A and B. Right Panel: Finally, we show the CDF of the in concordance $\log R$ distribution in the right panel of the figure and the corresponding CDF values for each pair of observations. We find that for the two in tension observations the $T=2.989^{+0.167}_{-0.060}$ and $T=2.147^{+0.056}_{-0.089}$ and for the in concordance observations $C=0.864^{+0.107}_{-0.076}$ and $T=0.507^{+0.063}_{-0.078}$. The results are in agreement with our expectations given the relative amplitudes of the observed signals, and the example demonstrates the application of the tensionnet on a problem with no analytically tractable in concordance $\log R$ distribution.
@@
We simulate observations of the sky-averaged 21-cm signal from two different experiments in different frequency ranges. We hypothesise that the 21-cm signal has a depth of 0.2 K, a central frequency of 78 MHz and a width of 10 MHz. In our example, the first experiment (Exp. A) has made a detection of the signal with Gaussian distributed noise with a standard deviation of 25 mK over the frequency range $60-90$ MHz with a channel width of $\approx 0.3$ MHz (see top left panel of fig. 5). We then hypothesise a series of scenarios where a second experiment (Exp. B) has observed the 21-cm signal in the frequency range $80-120$ MHz with a channel width of $\approx 0.4$ MHz with the same central frequency and width but a different magnitude $A=[0.15,0.2,0.25]$ K such that the observations are in tension, concordance and tension respectively. We add 25 mK Gaussian random noise to the data from experiment B.
@@
We fit each pair of observations using the nested sampling implementation polychord (Handley et al., 2015a, b) to assess $R_{\mathrm{obs}}$. We use eq. 21 as our model, $M$ for the data, $D$ and use a Gaussian likelihood function
@@
where the sum is over observation frequency and $\sigma$ is the standard deviation of the noise, which we fit as a free parameter. The prior is uniform on $A$ between $0.0-4.0$ K, $\nu_{0}$ between $60-80$ MHz, $w$ between $5-40$ MHz and $\sigma$ between $0.001-0.1$ K. The combination of our likelihood and prior and the fact that our model is non-linear makes the in concordance $\log R$ distribution analytically intractable. It can only be accessed in a reasonable amount of time through the tensionnet. One could of course evaluate the distribution with 1000s of Nested Sampling runs, but this would be computationally expensive.
@@
Once trained, we generate 5000 pairs of observations of the same signal by both experiment with parameters drawn randomly from the prior range to evaluate the in concordance $\log R$ distribution. From this distribution, we can calculate an empirical CDF and compare the observed $R$ statistic for the three pairs of observations. Nested sampling returns an error on the Bayesian evidence, which can then be propagated forward through to $\log R_{\mathrm{obs}}$ and the tension statistics $T$ and $C$. For the two in tension datasets we find $T=2.989^{+0.167}_{-0.060}$ and $T=2.147^{+0.056}_{-0.089}$ and for the in concordance case when both experiments observe the same signal $C=0.864^{+0.107}_{-0.076}$ and $T=0.507^{+0.063}_{-0.078}$. This is in agreement with our expectations, given the amplitude of the signals in the different data sets, and demonstrates that the tensionnet performs well. The results are summarised in fig. 5.
@@ VII.2 DESI and SDSS
We next investigate the tension between the Baryon Acoustic Oscillations (BAO) cosmological constraints from the Sloan Digital Sky Survey (SDSS) Alam et al. (2015); Ahumada et al. (2020) and the recent Dark Energy Spectroscopic Instrument (DESI) data release Adame et al. (2024).
Before recombination when photons and baryons were coupled via Thomson scattering, oscillations were set up in the hot plasma by the competing forces of gravity and radiation pressure. Spherical density perturbations in the coupled plasma propagated outwards as acoustic waves. Once the photons and the baryons decouple at recombination, these acoustic waves stop travelling through the baryon fluid and the scale of the wave is imprinted in the matter distribution. The scale of the acoustic waves at recombination is known as the sound horizon. The photons free stream and form the CMB. Since the baryons and dark matter are coupled by gravity, the acoustic waves imprint a preferential scale for structure formation and the distance between two galaxies in the later Universe. The BAO scale is hence a standard ruler, and observations of it can be used to constrain the expansion rate of and the matter density of the Universe Bassett and Hlozek (2010); Cuceu et al. (2019).
@@
The BAO signature appears in the cross-correlation of a number of different objects such as Luminous Red Galaxies (LRG), Emission Line Galaxies (ELG), quasars and the Lyman-$\alpha$ forest. Each class of objects probes a different redshift range, and measurements of $D_{M}/r_{s}$ and $D_{H}/r_{s}$ are ascribed to an effective redshift Bassett and Hlozek (2010).
We generate theoretical models for the observables with CAMB (Lewis et al., 2000; Lewis and Bridle, 2002), then taking advantage of the reported covariance estimates for the SDSS and DESI observations use analytic likelihoods, implemented with scipy, to generate noisy observations of the theory model.
There is a partial overlap in the redshift range and sky coverage of SDSS and DESI, and as such the full datasets include some of the same galaxies. Therefore, the surveys are correlated and if we want to perform a joint Bayesian inference of the datasets with tools like nested sampling to recover $R_{\mathrm{obs}}$ we need a joint likelihood function. Although the level of correlation between the datasets has been estimated DESI Collaboration (2024a); Adame et al. (2024), the derivation of a joint likelihood is beyond the scope of this paper and has not yet been attempted in the literature.
An alternative approach is to build a joint SDSS and DESI dataset by selecting data points from one survey or the other at each sampled effective redshift. In Adame et al. (2024) the authors demonstrate this idea by selecting SDSS observations below $z=0.6$ and DESI observations above $z=0.6$ to maximise the effective volume covered by the joint dataset. In our analysis we use
@@
Figure 6: The composite BAO dataset used in this work from SDSS and DESI observations. Following the discussion in Adame et al. (2024) we curate the dataset by taking SDSS observations of the BAO scale from Luminous Red Galaxies (LRG, crosses) below $z=0.6$ (grey shaded region) and high redshift observations of Luminous Red Galaxies and Emission Line Galaxies (ELG; combination of ELG and LRG as circles and ELG on their own as triangle markers) from DESI (blue shaded region). The measurements of $D_{M}/r_{s}$ are shown by the orange markers and the measurements of $D_{H}/r_{s}$ as purple markers.
@@
We constrain the baryon density $\Omega_{b}h^{2}$, dark matter density $\Omega_{c}h^{2}$, the slope and amplitude of the matter power spectrum $n_{s}$ and $\log 10^{10}A_{s}$ and the value of $h=\frac{H_{0}}{100~{}{\rm km~{}s}^{-1}~{}{\rm Mpc}^{-1}}$. We fix the value of $\tau$ to the best fit value from the Planck 2018 analysis of $0.055$ Planck Collaboration (2020). The prior is uniform on $\Omega_{b}h^{2}$ between $0.01-0.085$, $\Omega_{c}h^{2}$ between $0.08-0.21$, $n_{s}$ between $0.8-1.2$, $\log 10^{10}A_{s}$ between $2.6-3.8$ and $h$ between $0.5-0.9$. It is motivated by the prior in Handley and Lemos (2019), which is somewhat motivated by the default priors for CosmoMC  [^2] https://cosmologist.info/cosmomc/, and designed to encompass the Planck and Dark Energy Survey Y1 posteriors. As discussed in Handley and Lemos (2019), however, there is nothing particularly special about this prior and in practice it could be broadened or narrowed without causing any objections in the community. For each measurement of the BAO signature $D$ our likelihood is Gaussian, as in Cuceu et al. (2019), with a covariance given by the measured covariance $\Sigma$.
@@
The SDSS data is available at https://www.sdss4.org/dr17/ and the DESI data has been reported in Adame et al. (2024). Both datasets have been collected together as part of the COBAYA cosmological likelihood code  [^3] https://cobaya.readthedocs.io/en/latest/likelihood_bao.html. Using nested sampling and CAMB, we find $\log R_{\mathrm{obs}}=2.57\pm 0.30$. Since $R_{\rm obs}<10$ we are not worried about the NRE saturation that was previously discussed.
@@
To train the NRE, we generate 100,000 examples of in concordance observations from SDSS and DESI. We then separate out 10% of these for testing and shuffle the remaining $90\%$ to create a set of 180,000 matched and mismatched observations. These are then split into training and validation datasets of 120,600 and 59,400 (33%) observations respectively. We use an exponentially decaying learning rate scheduler with an initial learning rate of $10^{-3}$, a step size of $1000$ and a decay rate of $0.9$. We train for a maximum of 1000 epochs with a batch size of 1000 and a patience of 50. We use L1 kernel regularization to improve the performance. We standardize the simulations at each redshift using the mean and standard deviation of the training data.
@@
We group together the measurements of $D_{M}/r_{s}$ from DESI and SDSS at the different effective redshifts and compress them down into a smaller latent space. We do the same with the measurements of $D_{H}/r_{s}$ before passing them to the NRE. We find that compressing the data in this way works better than directly passing the raw data to the NRE. This initial step of keeping the measurements of $D_{M}/r_{s}$ and $D_{H}/r_{s}$ separate allows the NRE to learn the trends, like those seen in fig. 6, in each variable as a function of redshift before mixing information from the two together. The compression networks have three layers of 5, 5 and 2 hidden nodes and the NRE has 2 layers of 4 nodes each. The compression layers and the NRE are trained together. The architecture of the network can be seen in fig. 7.
@@
We train the network on the same training data five times with different random initial seeds to assess the consistency of our results. The corresponding values of $T$ are shown in fig. 8. We find that on average, $T=1.22\pm 0.20$ between the combined SDSS and DESI datasets. We show an example of the calibration performed for one of the training and calibration runs in fig. 9 along with the constraints on $\Omega_{m}$ and $H_{0}r_{s}$. We find no significant tension between the SDSS measurements of the BAO scale at $z_{\mathrm{eff}}=0.38$ and $0.51$ and the DESI measurements at higher redshifts of $z_{\mathrm{eff}}=0.706,0.930$ and $1.317$.
Figure 8: We repeat training of our NRE five times on simulations of the data from SDSS and DESI and use the predicted distributions to evaluate $T$. If our network was ill-converged, or we had two little training data, then the recovered distribution of in concordance $R$ values would vary significantly. As a result, the calculated value of $T$ would be inconsistent, and we would see a large scatter in the reported values. Instead, we see that for the SDSS+DESI analysis, $T$ is consistent across the different training runs. On average, we find that $T=1.22\pm 0.20$.
Figure 9: Left Panel: The predicted distribution of in concordance $\log R$ values for the curated SDSS and DESI dataset analysed in this work. The red dashed line shows the value of $\log R_{\mathrm{obs}}$ for the observed data calculated with nested sampling. The shaded region shows the error on this value from the nested sampling algorithm. Middle Panel: The CDF corresponding to the in concordance $\log R$ distribution. Calibrating $\log R_{\mathrm{obs}}$ (red dashed line and shaded region) in to $\sigma$s of tension gives $T=1.23^{+0.21}_{-0.20}$. Right Panel: The constraints on the matter overdensity $\Omega_{m}$ and the combination of the Hubble constant $H_{0}$ and sound horizon $r_{s}$ from analysis of the DESI and SDSS datasets. Here, the SDSS data comprises observations $z<0.6$ and the DESI data observations with $z>0.6$. Our results are slightly different to those presented in Fig. 2 of Adame et al. (2024) because we have used a different prior and not included the quasar measurements from DESI or the Lyman-$\alpha$ measurements from both surveys.

### v2

@@ VII.1 Toy 21-cm Cosmology
Observers in the field of 21-cm Cosmology are aiming to detect an information rich redshifted signal from neutral hydrogen from the Cosmic Dawn and Epoch of Reionization (see Furlanetto2006; Barkana2016; Mesinger2019; Liu2020, for reviews of the field). The signal is observed in the radio band, and can in theory be detected as a sky-averaged 21-cm signal (e.g. EDGES; SARAS3; Acedo2022REACH). It has a complex dependence on the astrophysics of the early Universe (e.g. Mirocha2014ARES; Mesinger201121cmfast; Reis2020radioGalaxies; Reis2021Lyalpha; GesseyJones2022IMF; Skider2024LineSight; Pochinda2023Joint; GesseyJones2024Joint; Munoz2023Zeus; Mittal2025Echo21) and can be thought of as a spectral distortion in the CMB. The key challenge in 21-cm cosmology is the separation of this signal from the dominant Galactic and extragalactic foregrounds, that the instruments also observe, whilst accounting for the non-uniform response of the instruments to the sky (Anstey2021REACH).
Since we are focused on illustrating the performance of the tensionnet, we ignore the effects of foregrounds and the instrument in our example and use a simplistic gaussian absorption feature to model the 21-cm signal. We include Gaussian distributed noise in our simulated data (inspired by current observations EDGES; SARAS3) and a Gaussian absorption feature
@@
where $A$ corresponds to the amplitude of the signal, $\nu_{0}$ to the central frequency and $w$ to the width. A number of works in the literature have adopted this model for illustrative purposes (e.g Anstey2021REACH; Bevins2021maxsmooth) but we stress that it is simplistic and does not capture the complex astrophysics encoded by the 21-cm signal.
@@
Current observations of the sky-averaged 21-cm signal include a tentative detection by the EDGES collaboration EDGES and an upper limit on the magnitude of the signal from SARAS3 SARAS3. Analysis by the SARAS3 team suggested that these measurements are in tension with each other, and a number of works have discussed the possible presence of systematics in the EDGES data Hills2018Edges; Singh2019EDGES; Sims2020EDGES; Bevins2021maxsmooth. As more experiments come online in the coming years (e.g. Acedo2022REACH) the assessment of tension and concordance between different observations is going to become crucial for the field.
@@
Figure 5: To further illustrate the application of NREs to the calibration of $R$ we use a toy example inspired by 21-cm cosmology. Left Panel: We simulate an experiment observing a Gaussian absorption trough as a function of frequency (black line) and three different scenarios in which another experiment measures a 21-cm signal with either the same or different amplitudes in a different band (red lines). To each observation, we add Gaussian random noise with a standard deviation of 25 mK (shown in grey and motivated by current observations (e.g. EDGES)). Middle Panel: We train the NRE on simulated observations of the signal by both experiments, covering a wide prior range of signal parameters. We use the NRE to evaluate the possible distribution of in concordance $\log R$ values. We plot the observed $\log R$ for each pair of observations from experiment A and B. Right Panel: Finally, we show the CDF of the in concordance $\log R$ distribution in the right panel of the figure and the corresponding CDF values for each pair of observations. We find that for the two in tension observations the $T=2.989^{+0.167}_{-0.060}$ and $T=2.147^{+0.056}_{-0.089}$ and for the in concordance observations $C=0.864^{+0.107}_{-0.076}$ and $T=0.507^{+0.063}_{-0.078}$.
@@
We simulate observations of the sky-averaged 21-cm signal from two different experiments in different frequency ranges. We hypothesise that the 21-cm signal has a depth of 0.2 K, a central frequency of 78 MHz and a width of 10 MHz. In our example, the first experiment (Exp. A) has made a detection of the signal with Gaussian distributed noise with a standard deviation of 25 mK over the frequency range 609060-90 MHz with a channel width of $\approx 0.3$ MHz (see top left panel of fig.˜5). We then hypothesise a series of scenarios where a second experiment (Exp. B) has observed the 21-cm signal in the frequency range 8012080-120 MHz with a channel width of $\approx 0.4$ MHz with the same central frequency and width but a different magnitude $A=[0.15,0.2,0.25]$ K such that the observations are in tension, concordance and tension respectively. We add 25 mK Gaussian random noise to the data from experiment B.
@@
We fit each pair of observations using the nested sampling implementation polychord (Handley2015polychorda; Handley2015polychordb) to assess $R_{\mathrm{obs}}$. We use eq.˜21 as our model, $M$ for the data, $D$ and use a Gaussian likelihood function
@@
where the sum is over observation frequency and $\sigma$ is the standard deviation of the noise, which we fit as a free parameter. The prior is uniform on $A$ between 0.04.00.0-4.0 K, $\nu_{0}$ between 608060-80 MHz, $w$ between 5405-40 MHz and $\sigma$ between 0.0010.10.001-0.1 K.
The combination of our likelihood and prior and the fact that our model is non-linear makes the in concordance $\log R$ distribution analytically intractable. It can only be accessed in a reasonable amount of time through the tensionnet. One could of course evaluate the distribution with 1000s of Nested Sampling runs, but this would be computationally expensive. To calculate $R$ using Nested Sampling for this example costs around 12 seconds (all the analysis performed in the paper was done on an M2 MacBook with 8GB ram). This includes the run time of the inference on each data set individually and on the joint data set. Using Nested Sampling to evaluate $R$ for 5000 pairs of simulated in concordance, observations to assess the distribution of possible values, would therefore take around 16.7 hrs. To generate simulations for the NRE, train the NRE and evaluate it for 5000 pairs of data sets to get the in concordance $R$ distribution takes around 6.6 minutes. For this problem, the NRE is therefore $(16.7\times 60)/6.6\approx 152$ times faster. Note that this is a trivial problem with a fast likelihood function.
@@
Once trained, we generate 5000 pairs of observations of the same signal by both experiment with parameters drawn randomly from the prior range to evaluate the in concordance $\log R$ distribution. From this distribution, we can calculate an empirical CDF and compare the observed $R$ statistic for the three pairs of observations. Nested sampling returns an error on the Bayesian evidence, which can then be propagated forward through to $\log R_{\mathrm{obs}}$ and the tension statistics $T$ and $C$. For the two in tension datasets we find $T=2.989^{+0.167}_{-0.060}$ and $T=2.147^{+0.056}_{-0.089}$ and for the in concordance case when both experiments observe the same signal $C=0.864^{+0.107}_{-0.076}$ and $T=0.507^{+0.063}_{-0.078}$. This is in agreement with our expectations, given the amplitude of the signals in the different data sets, and demonstrates that the tensionnet performs well. The results are summarised in fig.˜5.
@@ VII.2 DESI and SDSS
We next investigate the tension between the Baryon Acoustic Oscillations (BAO) cosmological constraints from the Sloan Digital Sky Survey (SDSS) Alam2015SDSSDR12; Ahumada2020SDSSDR16 and the recent Dark Energy Spectroscopic Instrument (DESI) data release adame2024desiCosmologyBAO.
Before recombination when photons and baryons were coupled via Thomson scattering, oscillations were set up in the hot plasma by the competing forces of gravity and radiation pressure. Spherical density perturbations in the coupled plasma propagated outwards as acoustic waves. Once the photons and the baryons decouple at recombination, these acoustic waves stop travelling through the baryon fluid and the scale of the wave is imprinted in the matter distribution. The scale of the acoustic waves at recombination is known as the sound horizon. The photons free stream and form the CMB. Since the baryons and dark matter are coupled by gravity, the acoustic waves imprint a preferential scale for structure formation and the distance between two galaxies in the later Universe. The BAO scale is hence a standard ruler, and observations of it can be used to constrain the expansion rate of and the matter density of the Universe Bassett2010BAO; Cuceu2019BAO.
@@
The BAO signature appears in the cross-correlation of a number of different objects such as Luminous Red Galaxies (LRG), Emission Line Galaxies (ELG), quasars and the Lyman-$\alpha$ forest. Each class of objects probes a different redshift range, and measurements of $D_{M}/r_{s}$ and $D_{H}/r_{s}$ are ascribed to an effective redshift Bassett2010BAO.
We generate theoretical models for the observables with CAMB (Lewis1999CAMB; Lewis2002CAMB), then taking advantage of the reported covariance estimates for the SDSS and DESI observations use analytic likelihoods, implemented with scipy, to generate noisy observations of the theory model.
There is a partial overlap in the redshift range and sky coverage of SDSS and DESI, and as such the full datasets include some of the same galaxies. Therefore, the surveys are correlated and if we want to perform a joint Bayesian inference of the datasets with tools like nested sampling to recover $R_{\mathrm{obs}}$ we need a joint likelihood function. Although the level of correlation between the datasets has been estimated adame2024desiLyaBAO; adame2024desiCosmologyBAO, the derivation of a joint likelihood is beyond the scope of this paper and has not yet been attempted in the literature.
An alternative approach is to build a joint SDSS and DESI dataset by selecting data points from one survey or the other at each sampled effective redshift. In adame2024desiCosmologyBAO the authors demonstrate this idea by selecting SDSS observations below $z=0.6$ and DESI observations above $z=0.6$ to maximise the effective volume covered by the joint dataset. In our analysis we use
@@
Figure 6: The composite BAO dataset used in this work from SDSS and DESI observations. Following the discussion in adame2024desiCosmologyBAO we curate the dataset by taking SDSS observations of the BAO scale from Luminous Red Galaxies (LRG) below $z=0.6$ (grey shaded region) and high redshift observations of Luminous Red Galaxies and Emission Line Galaxies (ELG) from DESI (blue shaded region).
@@
We constrain the baryon density $\Omega_{b}h^{2}$, dark matter density $\Omega_{c}h^{2}$, the slope and amplitude of the matter power spectrum $n_{s}$ and $\log 10^{10}A_{s}$ and the value of $h=\frac{H_{0}}{100~{\rm km~s}^{-1}~{\rm Mpc}^{-1}}$. We fix the value of $\tau$ to the best fit value from the Planck 2018 analysis of $0.055$ Planck2018. The prior is uniform on $\Omega_{b}h^{2}$ between 0.010.0850.01-0.085, $\Omega_{c}h^{2}$ between 0.080.210.08-0.21, $n_{s}$ between 0.81.20.8-1.2, $\log 10^{10}A_{s}$ between 2.63.82.6-3.8 and $h$ between 0.50.90.5-0.9. It is motivated by the prior in Handley_tensions_2019, which is somewhat motivated by the default priors for CosmoMC  [^2] https://cosmologist.info/cosmomc/, and designed to encompass the Planck and Dark Energy Survey Y1 posteriors. As discussed in Handley_tensions_2019, however, there is nothing particularly special about this prior and in practice it could be broadened or narrowed without causing any objections in the community. For each measurement of the BAO signature $D$ our likelihood is Gaussian, as in Cuceu2019BAO, with a covariance given by the measured covariance $\Sigma$.
@@
The SDSS data is available at https://www.sdss4.org/dr17/ and the DESI data has been reported in adame2024desiCosmologyBAO. Both datasets have been collected together as part of the COBAYA cosmological likelihood code  [^3] https://cobaya.readthedocs.io/en/latest/likelihood_bao.html. Using nested sampling and CAMB, we find $\log R_{\mathrm{obs}}=2.57\pm 0.30$. Since $R_{\rm obs}<10$ we are not worried about the NRE saturation that was previously discussed.
@@
To train the NRE, we generate 100,000 examples of in concordance observations from SDSS and DESI. We then separate out 10% of these for testing and shuffle the remaining $90\%$ to create a set of 180,000 matched and mismatched observations. These are then split into training and validation datasets of 120,600 and 59,400 (33%) observations respectively. We use an exponentially decaying learning rate scheduler with an initial learning rate of $10^{-3}$, a step size of $1000$ epochs and a decay rate of $0.9$. We train for a maximum of 1000 epochs with a batch size of 1000 and a patience of 50. We use L1 kernel regularization to improve the performance. We standardize the simulations at each redshift using the mean and standard deviation of the training data.
@@
We group together the measurements of $D_{M}/r_{s}$ from DESI and SDSS at the different effective redshifts and compress them down into a smaller latent space. We do the same with the measurements of $D_{H}/r_{s}$ before passing them to the NRE. We find that compressing the data in this way works better than directly passing the raw data to the NRE. This initial step of keeping the measurements of $D_{M}/r_{s}$ and $D_{H}/r_{s}$ separate allows the NRE to learn the trends, like those seen in fig.˜6, in each variable as a function of redshift before mixing information from the two together. We found that smaller neural network architectures worked better for this problem, as there are only 10 data points in the combined set. The compression networks have three layers of 5, 5 and 2 hidden nodes and the NRE has 2 layers of 4 nodes each. The compression layers and the NRE are trained together. The architecture of the network can be seen in fig.˜7.
$\log R$ SDSS LRG $z_{\mathrm{eff}}=0.38$ SDSS LRG $z_{\mathrm{eff}}=0.51$ DESI LRG $z_{\mathrm{eff}}=0.706$ DESI LRG-ELG $z_{\mathrm{eff}}=0.930$ DESI LRG $z_{\mathrm{eff}}=1.317$ SDSS LRG $z_{\mathrm{eff}}=0.38$ SDSS LRG $z_{\mathrm{eff}}=0.51$ DESI LRG $z_{\mathrm{eff}}=0.706$ DESI LRG-ELG $z_{\mathrm{eff}}=0.930$ DESI LRG $z_{\mathrm{eff}}=1.317$ $D_{M}/r_{s}$ { $D_{H}/r_{s}$ {
We train the network on the same training data five times with different random initial seeds to assess the consistency of our results. The corresponding values of $T$ are shown in fig.˜8. We find that on average, $T=1.22\pm 0.20$ between the combined SDSS and DESI datasets. We show an example of the calibration performed for one of the training and calibration runs in fig.˜9 along with the constraints on $\Omega_{m}$ and $H_{0}r_{s}$. We find no significant tension between the SDSS measurements of the BAO scale at $z_{\mathrm{eff}}=0.38$ and $0.51$ and the DESI measurements at higher redshifts of $z_{\mathrm{eff}}=0.706,0.930$ and $1.317$.
Figure 8: We repeat training of our NRE five times on simulations of the data from SDSS and DESI and use the predicted distributions to evaluate $T$. If our network was poorly converged, or we had too little training data, then the recovered distribution of in concordance $R$ values would vary significantly. As a result, the calculated value of $T$ would be inconsistent, and we would see a large scatter in the reported values. Instead, we see that, for the SDSS+DESI analysis, $T$ is consistent across the different training runs. On average, we find that $T=1.22\pm 0.20$.